# Maximum Implementable Improvements over Deferred Acceptance

### v2

@@ 4 Limitations of Existing Mechanisms
How many students do existing mechanisms like EADA and DA+TTC actually improve compared to the maximum improvement possible over DA? To quantify this gap, we define the improvement ratio of a mechanism as follows. Let $\mathcal{I}(M,P)$ denote the set of students improved by the mechanism $M$ that Pareto-dominates DA in problem $P$:
$\mathcal{I}(M,P)\coloneqq\{i\in I:\textup{rk}_{i}(M_{i}(P))<\textup{rk}_{i}(\mathtt{DA}_{i}(P))\}$ (2)
We define the maximum improvement $\mathcal{I}(\mu^{*},P)$ as:
$\mathcal{I}(\mu^{*},P)\coloneqq\max_{\mu\in\mathcal{M}^{\mathtt{DA}}(P)}|\mathcal{I}(\mu,P)|$ (3)
@@
$\rho(M;n)\coloneqq\max_{P\in\mathcal{P}_{n}}\frac{|\mathcal{I}(\mu^{*},P)|}{|\mathcal{I}(M,P)|}$ (4)
For example, an improvement ratio of 2 means that the maximum mechanism may generate twice as many improvements as mechanism $M$. Our first result shows that this ratio can grow linearly with the size of the problem. [^4] Throughout Sections 4-4.1, we assume all students consent to waive their priorities, as even with full consent we obtain the worst possible bounds. We examine partial consent explicitly in Section 4.2 and formalize arbitrary consent structures in Section 5.
@@
For every $n>4$ there exists a problem $P$ such that the maximum improvement is $n-1$; while that of EADA and DA+TTC are both exactly 2:
$\rho(\text{EADA};n)=\rho(\text{DA+TTC};n)=\frac{n-1}{2}$
@@
Upper Bound. First we prove that the improvement ratio cannot exceed $\frac{n-1}{2}$. For the numerator: At most $n-1$ can improve over DA, as there is always at least one unimprovable student (Tang and Yu, 2014). For the denominator: The minimum number of students improved by EADA and DA+TTC is 2, because a trading cycle requires at least two students. [^5] We have assumed that DA is inefficient; otherwise, the minimum would be 0 and the ratio undefined. Therefore, the ratio cannot exceed $\frac{n-1}{2}$.
Tightness. We construct a school choice problem with $n$ students and schools (all with unit capacity) where this improvement ratio is achieved for an arbitrary value of $n>4$.
@@
The table below demonstrates this construction for $n=7$. The DA allocation appears in bold. The allocation that maximises the number of improvements appears in circles, whereas the matching recommended by both EADA and DA+TTC appears in squares (whenever different from DA).
@@
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\bm{s_{5}}$ $\bm{s_{6}}$cdot $i_{2}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
@@
Analysis of EADA.
Under EADA, we begin by identifying under-demanded schools. First, $s_{n}$ is under-demanded and is removed together with $i_{n}$. In the next iteration, $s_{3},s_{4},\ldots,s_{n-1}$ are under-demanded and removed jointly with $i_{3},i_{4},\ldots,i_{n-1}$, respectively. At this stage, only schools $s_{1}$ and $s_{2}$ remain with students $i_{1}$ and $i_{2}$. These students form a trading cycle where $i_{1}$ improves to $s_{2}$ and $i_{2}$ improves to $s_{1}$. Thus, EADA improves exactly 2 students (assuming $i_{n}$ consents to waive his priority at $s_{1}$, otherwise no student improves).
@@
There exists, however, a matching $\mu^{*}$ that dominates DA, obtained by implementing the cycles $(i_{1}\rightarrow i_{n-1}\rightarrow i_{1})$ and $(i_{2}\rightarrow i_{3}\rightarrow\ldots\rightarrow i_{n-2}\rightarrow i_{2})$, which improves $n-1$ students. Under this matching, student $i_{n}$ receives the same allocation as under DA, while all other students receive a more preferred school.
@@
Therefore, the improvement ratio equals $\frac{|I(\mu^{*},P)|}{|I(M,P)|}=\frac{n-1}{2}$ for any $n>4$, establishing the tightness of our bound. ∎
@@
Three remarks are in order now. First, Kesten (2010) already pointed out in his seminal paper that larger improvements than EADA were possible. Our contribution here is to quantify how much larger those improvements can be, leading to an improvement gap that grows linearly with the problem size.
Second, in our constructed example, the maximum improvement also dominates on rank positions improved: the 2-student cycle generates 4 total rank improvements while the $(n-1)$-student cycle generates $n$ total rank improvements. For $n\geq 6$, maximizing student count yields strictly higher rank improvement, showing that existing mechanisms can simultaneously underperform on both distributional and rank-efficiency dimensions, albeit this need not be the case in general.
@@
Third, in our example, the maximum improvement comes at the cost of generating more blocking pairs. In the particular example with seven agents, the maximum improvement generates at least three blocking pairs ($\{i_{1},s_{2}\},\{i_{2},s_{1}\},\{i_{7},s_{1}\}$) versus a single blocking pair generated by EADA and DA+TTC ($\{i_{7},s_{1}\}$). We do not take a stance on whether the larger improvement is worth the additional justified envy, our point is to make explicit that such reduction in justified envy comes at a high cost in terms of how many students are benefited. Furthermore, as we will show next, the larger improvement can sometimes lead to fewer blocking pairs.
@@ 4.1 Larger Improvements with Less Justified Envy
Theorem 1 shows that EADA and DA+TTC can improve significantly fewer students, establishing a tight bound. In the example used in the proof, we showed that larger improvements came at a cost of more blocking pairs. Now we will show that this need not always be the case: in some problems, EADA and DA+TTC can be doubly dominated: they may improve fewer students while generating more justified envy.
Proposition 1.
There exist school choice problems where a larger improvement than EADA’s and DA+TTC’s is possible that generates fewer blocking pairs.
@@
Comparison with EADA. Consider the following school choice problem with seven students and seven schools with one seat each. DA’s allocation appears in bold. Preferences and priorities left unspecified are irrelevant.
@@
$s_{6}$ $s_{1}$ $s_{6}$ $s_{5}$ $s_{3}$ $s_{4}$ $s_{4}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$cdot
$s_{4}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $s_{6}$ $\boldsymbol{s_{6}}$ $\boldsymbol{s_{7}}$ $i_{5}$ $i_{1}$ $i_{5}$ $i_{7}$ $i_{4}$ $i_{3}$cdot
$s_{2}$cdot $\cdot$ $\cdot$ $s_{4}$cdot $\cdot$ $i_{2}$cdot $i_{1}$ $i_{1}$ $i_{1}$ $i_{5}$cdot
$s_{3}$cdot $\cdot$ $\cdot$ $s_{1}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$cdot $i_{1}$cdot
$s_{5}$cdot $\cdot$ $\cdot$ $\bm{s_{5}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{5}$cdot $\cdot$ $\cdot$
$\bm{s_{1}}$cdot $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
The DA envy digraph appears below, showing the trading cycles implemented by EADA (Figure 1, improving 4 students through the cycle $i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1}$) and the maximum improvement (Figure 2, improving 6 students via two disjoint cycles, namely $i_{1}\to i_{2}\to i_{1}$ and $i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3}$). Interestingly, EADA generates 3 blocking pairs ($\{i_{3},s_{6}\},\{i_{5},s_{6}\},\{i_{7},s_{4}\}$), whereas the maximum improvement generates only two blocking pairs ($\{i_{1},s_{4}\},\{i_{7},s_{4}\}$). [^6] Note that here we are simply counting the number of blocking pairs; there is a literature (see Section 2) arguing that that some are more relevant than others. Note that the maximum improvement improves every improvable student, because $i_{7}$ is unimprovable since nobody envies him.
@@
Figure 1: EADA
@@
Figure 2: Maximum Improvement
Comparison with DA+TTC. To show that a larger improvement can generate fewer blocking pairs than DA+TTC and can be implemented with fewer students consenting, consider the following example with five students and five schools with unit capacity. DA’s allocation appears in bold.
@@
$s_{4}$ $s_{3}$ $s_{2}$ $s_{2}$ $\bm{s_{5}}$ $i_{5}$ $i_{4}$ $i_{4}$ $i_{5}$cdot
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$cdot $i_{4}$ $i_{3}$ $i_{2}$ $i_{1}$cdot
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{2}$ $i_{1}$cdot $\cdot$ $\cdot$
In DA+TTC, students $i_{1}$ and $i_{2}$ trade their DA allocations, generating four blocking pairs: $(i_{4},s_{1}),(i_{5},s_{1}),(i_{3},s_{2}),(i_{4},s_{2})$. In contrast, the maximum improvement is the cycle $(i_{1}\to i_{4}\to i_{1}),(i_{2}\to i_{3}\to i_{2})$, generating only one blocking pair: $(i_{5},s_{1})$. ∎
While DA+TTC, and in particular, EADA, choose improvement cycles taking priorities into account, therefore satisfying several weak stability properties (see Section 2), Proposition 1 shows that they may generate doubly dominated matchings in some problems—improving fewer students while creating more blocking pairs—since alternative matchings can achieve better outcomes on both dimensions simultaneously.
Noting that the larger improvements in the previous two examples are obtained by DA+TTC and EADA, respectively, yields the following incomparability corollary.
@@
EADA and DA+TTC are incomparable in terms of size of improvement and number of blocking pairs, i.e. there exist problems where DA+TTC doubly dominates EADA and vice versa.

### v3

@@ 4 Limitations of Existing Mechanisms
How many students do existing mechanisms actually improve compared to the maximum improvement possible over DA? To quantify this gap, we define the improvement ratio of a mechanism. Let $\mathcal{I}(\mu,P)$ denote the set of students improved by the matching $\mu$ that Pareto-dominates DA in problem $P$:
$\mathcal{I}(\mu,P)\coloneqq\{i\in I:\textup{rk}_{i}(\mu_{i})<\textup{rk}_{i}(\mathtt{DA}_{i}(P))\}$ (2)
Let $\mathcal{M}^{\mathtt{DA}}(P)$ denote the set of matchings that weakly Pareto-dominate DA in $P$.
We define the maximum improvement $\mathcal{I^{*}}(P)$ as:
$\mathcal{I^{*}}(P)\coloneqq\max_{\mu\in\mathcal{M}^{\mathtt{DA}}(P)}|\mathcal{I}(\mu,P)|$ (3)
@@
$\rho(M;n)\coloneqq\max_{P\in\mathcal{P}_{n}}\frac{|\mathcal{I^{*}}(P)|}{|\mathcal{I}(M(P),P)|}$ (4)
For example, an improvement ratio of 2 means that the mechanism that improves as many students as possible compared to DA may generate twice as many improvements as mechanism $M$. Our first result shows that this ratio can grow linearly with the size of the problem, and particularly that the improvement ratio of DA+TTC and EADA with full consent achieve the worst possible improvement ratio among all mechanisms that weakly Pareto-dominate DA.
@@
For any $n>4$ and any Pareto-efficient mechanism $M$ that Pareto-dominates DA:
$\rho(\mathtt{EADA};n)=\rho(\mathtt{DA+TTC};n)=\frac{n-1}{2}\geq\rho(M;n)$
@@
Upper Bound. First we prove that the improvement ratio cannot exceed $\frac{n-1}{2}$ for any mechanism that Pareto-dominates DA. For the numerator: At most n1n-1 students can improve over DA, as there is always at least one unimprovable student (Tang and Yu, 2014). For the denominator: The minimum number of students improved by any Pareto-efficient mechanism that Pareto-dominates DA is 2, because a trading cycle requires at least two students. [^4] We have assumed that DA is inefficient; otherwise, the minimum would be 0 and the ratio undefined. Therefore, the ratio cannot exceed $\frac{n-1}{2}$.
Tightness. To show this bound is achievable, we construct a school choice problem with $n$ students and schools (all with unit capacity) where EADA with full consent and DA+TTC achieve this improvement ratio while a different matching improves n1n-1 students.
@@
The table below demonstrates this construction for $n=7$. The DA allocation appears in bold. The allocation that maximises the number of improvements appears in circles, whereas the matching recommended by both EADA with full consent and DA+TTC appears in squares (whenever different from DA).
@@
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\bm{s_{5}}$ $\bm{s_{6}}$ $\cdot$ $i_{2}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
@@
Analysis of EADA with full consent.
We begin by identifying under-demanded schools. First, $s_{n}$ is under-demanded and is removed together with $i_{n}$. In the next iteration, $s_{3},s_{4},\ldots,s_{n-1}$ are under-demanded and removed jointly with $i_{3},i_{4},\ldots,i_{n-1}$, respectively. At this stage, only schools $s_{1}$ and $s_{2}$ remain with students $i_{1}$ and $i_{2}$. These students form a trading cycle where $i_{1}$ improves to $s_{2}$ and $i_{2}$ improves to $s_{1}$. Thus, EADA with full consent improves exactly 2 students (assuming $i_{n}$ consents to waive his priority at $s_{1}$, otherwise no student improves).
@@
There exists, however, a matching $\mu^{*}$ that dominates DA, obtained by implementing the cycles $(i_{1}\rightarrow i_{n-1}\rightarrow i_{1})$ and $(i_{2}\rightarrow i_{3}\rightarrow\ldots\rightarrow i_{n-2}\rightarrow i_{2})$, which improves n1n-1 students. Under this matching, student $i_{n}$ receives the same allocation as under DA, while all other students receive a more preferred school.
@@
Therefore, the improvement ratio equals $\frac{n-1}{2}$ for any $n>4$, establishing the tightness of our bound. ∎
@@
Some remarks are now in order. First, Kesten (2010) already pointed out in his seminal paper that larger improvements than EADA were possible. Our contribution here is to quantify how much larger those improvements can be even when all students consent to waive their priorities, leading to an improvement gap that grows linearly with the problem size.
@@
Second, in our example, the maximum improvement comes at the cost of generating more blocking pairs. In the particular example with seven agents, the larger improvement generates at least three blocking pairs ($\{i_{1},s_{2}\},\{i_{2},s_{1}\},\{i_{7},s_{1}\}$) versus a single blocking pair generated by EADA with full consent and DA+TTC ($\{i_{7},s_{1}\}$). We do not take a stance on whether the larger improvement is worth the additional justified envy. Our point is to make explicit that such reduction in justified envy comes at a high cost in terms of how many students are benefited. Furthermore, as we will show in the next subsection, larger improvements can sometimes lead to fewer blocking pairs.
Third, it is natural to ask whether EADA with full consent and DA+TTC are the only mechanism that fails to implement larger improvements that benefit more students. We answer this question in the negative in the next Subsection.
4.1 A General Impossibility Result
We now generalize Theorem 1 to a wide class of mechanisms beyond EADA with full consent and DA+TTC, namely all those that are setwise minimally unstable, i.e. those which never produce a set of blocking pairs that is strictly larger by set inclusion than that generated by another Pareto-efficient mechanism that Pareto-dominates DA.
Theorem 2.
Among Pareto-efficient mechanisms that Pareto-dominate DA, every setwise minimally unstable mechanism achieves an improvement ratio of $\frac{n-1}{2}$.
Proof.
We use the same problem as in the proof of Theorem 1. In the case with seven students and schools, there are six cycles in DA’s envy digraph, summarized in Table 1 below.
Table 1: Cycles and Blocking Pairs in Example 1
Cycles Blocking Pairs
$C_{1}=(i_{1}\to i_{2}\to i_{1})$ $(i_{7},s_{1})$
$C_{2}=(i_{1}\to i_{2}\to i_{3}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$C_{3}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$C_{4}=(i_{1}\to i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{1})$ $(i_{7},s_{1}),(i_{6},s_{1}),(i_{2},s_{1})$
$C_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ $(i_{1},s_{2})$
$C_{6}=(i_{1}\to i_{6}\to i_{1})$ $(i_{7},s_{1}),(i_{2},s_{1}),(i_{5},s_{6})$
$C_{7}=(i_{1}\to i_{6}\to i_{2}\to i_{1})$ $(i_{7},s_{1}),(i_{1},s_{2}),(i_{5},s_{6})$
Any Pareto-efficient mechanism that Pareto-dominates DA must implement at least one of these cycles (potentially two if mutually disjoint). In particular, the larger improvement that includes as many students as possible implements $C_{5}=(i_{2}\to i_{3}\to i_{4}\to i_{5}\to i_{2})$ and $C_{6}=(i_{1}\to i_{6}\to i_{1})$.
The cycles $C_{2},C_{3},C_{4},C_{6},C_{7}$ all generate a set of blocking pairs that is strictly larger (by set inclusion) than that generated by $C_{1}$. The cycle $C_{5}$ is incomparable to $C_{1}$, but implementing it by itself does not lead to a Pareto-efficient allocation, since the cycle $C_{6}$ could be jointly implemented. But if both $C_{5}$ and $C_{6}$ are implemented, the set of blocking pairs becomes $\{(i_{1},s_{2}),(i_{2},s_{1}),(i_{7},s_{1})\}$, a strict superset of the set of blocking pairs generated by $C_{1}$. Thus, we conclude that every setwise minimally unstable must implement $C_{1}$, therefore missing the larger improvements dictated by $C_{5}$ and $C_{6}$ jointly, leading to an improvement ratio of $\frac{n-1}{2}=\frac{6}{2}$.
Finally, the extension to arbitrary $n$ is straightforward: there are $n$ cycles, with $C_{k}=(i_{1}\to\ldots\to i_{k+1}\to i_{1})$ for all integer $1\leq k\leq n-3$, all of them generating the blocking pair ($i_{n},s_{1}$), and only $C_{1}$ introducing exactly this blocking pair. There are exactly three more cycles: $C_{n-2}=(i_{2}\to i_{3}\to\ldots\to i_{n-3}\to i_{2})$, $C_{n-1}=(i_{1}\to i_{6}\to i_{1})$, and $C_{n}=(i_{1}\to i_{n-1}\to i_{2}\to i_{1})$. Only one collection of cycles that leaves the graph acyclic is setwise minimally unstable: $C_{1}$. However, the cycles $C_{n-2}$ and $C_{n-1}$ are disjoint (and thus compatible) and lead to an allocation where n1n-1 students improve, and therefore any setwise minimally unstable mechanism incurs in the largest improvement ratio possible, namely $\frac{n-1}{2}$. ∎
Noting that EADA is setwise minimally unstable (as proven by Tang and Zhang (2021) and Doğan and Ehlers (2021a)) makes it evident that Theorem 2 implies Theorem 1. Nonetheless, we present both Theorems independently to simplify the argument and improve their readability.
@@ 4.2 Larger Improvements with Less Justified Envy
Theorems 1 and 2 shows that EADA with full consent and DA+TTC can improve significantly fewer students compared to the DA baseline, but that such smaller coverage is unavoidable if we insist on mechanisms that are setwise minimally unstable. Now we show that the impossibility disappears if we compare blocking pairs simply by counting them, i.e. we consider cardinal minimal instability.
To be precise, let $M$ and $N$ be Pareto-efficient mechanisms that Pareto-dominate DA. We say that $M$ improves more students than $N$ if for every problem $P$, we have $|\mathcal{I}(N(P),P)|\leq|\mathcal{I}(M(P),P)|$, with strict inequality for some $P$. We are now ready to state our result.
Theorem 3.
There exists a Pareto-efficient mechanism that Pareto-dominates DA that improves more students than EADA with full consent while being cardinally more stable than EADA with full consent. An analogous result holds for DA+TTC.
@@
Comparison with EADA with full consent. Consider the following school choice problem with seven students and seven schools with one seat each. DA’s allocation appears in bold. Preferences and priorities left unspecified are irrelevant.
@@
$s_{6}$ $s_{1}$ $s_{6}$ $s_{5}$ $s_{3}$ $s_{4}$ $s_{4}$ $\bm{i_{1}}$ $\bm{i_{2}}$ $\bm{i_{3}}$ $\bm{i_{4}}$ $\bm{i_{5}}$ $\bm{i_{6}}$ $\cdot$
$s_{4}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $s_{6}$ $\boldsymbol{s_{6}}$ $\boldsymbol{s_{7}}$ $i_{5}$ $i_{1}$ $i_{5}$ $i_{7}$ $i_{4}$ $i_{3}$ $\cdot$
$s_{2}$ $\cdot$ $\cdot$ $\cdot$ $s_{4}$ $\cdot$ $\cdot$ $i_{2}$ $\cdot$ $i_{1}$ $i_{1}$ $i_{1}$ $i_{5}$ $\cdot$
$s_{3}$ $\cdot$ $\cdot$ $\cdot$ $s_{1}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{6}$ $\cdot$ $i_{1}$ $\cdot$
$s_{5}$ $\cdot$ $\cdot$ $\cdot$ $\bm{s_{5}}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{5}$ $\cdot$ $\cdot$ $\cdot$
$\bm{s_{1}}$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
The DA envy digraph appears below, showing the trading cycles implemented by EADA with full consent (Figure 1), improving 4 students through the cycle $(i_{1}\to i_{6}\to i_{4}\to i_{5}\to i_{1})$ and an alternative, larger improvement (Figure 2), benefiting 6 students via two disjoint cycles, namely $(i_{1}\to i_{2}\to i_{1})$ and $(i_{3}\to i_{6}\to i_{4}\to i_{5}\to i_{3})$. Interestingly, EADA with full consent generates 3 blocking pairs ($\{i_{3},s_{6}\},\{i_{5},s_{6}\},\{i_{7},s_{4}\}$), whereas the alternative, larger improvement generates only two blocking pairs ($\{i_{1},s_{4}\},\{i_{7},s_{4}\}$). [^5] Note that here we are simply counting the number of blocking pairs; there is a literature (see Section 2) arguing that that some are more relevant than others. Note that the larger improvement benefits every student who could benefit under some allocation that Pareto-dominates DA, because $i_{7}$ cannot trade with anyone since no student envies him.
@@
Figure 1: EADA, full consent.
@@
Figure 2: Larger improvement.
To conclude the proof, consider the mechanism $M$ that chooses the same matching as EADA with full consent in every problem except the aforementioned one. Such matching clearly is Pareto-efficient, Pareto-dominates DA, improves more students than EADA with full consent and is cardinally more stable than EADA.
Comparison with DA+TTC. To show that a larger improvement can generate fewer blocking pairs than DA+TTC, consider the following example with five students and five schools with unit capacity. DA’s allocation appears in bold.
@@
$s_{4}$ $s_{3}$ $s_{2}$ $s_{2}$ $\bm{s_{5}}$ $i_{5}$ $i_{4}$ $i_{4}$ $i_{5}$ $\cdot$
$\bm{s_{1}}$ $\bm{s_{2}}$ $\bm{s_{3}}$ $\bm{s_{4}}$ $\cdot$ $i_{4}$ $i_{3}$ $i_{2}$ $i_{1}$ $\cdot$
$\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $i_{2}$ $i_{1}$ $\cdot$ $\cdot$ $\cdot$
In DA+TTC, students $i_{1}$ and $i_{2}$ trade their DA allocations, generating four blocking pairs: $(i_{4},s_{1}),(i_{5},s_{1}),(i_{3},s_{2}),(i_{4},s_{2})$. In contrast, the largest improvement is the cycle $(i_{1}\to i_{4}\to i_{1}),(i_{2}\to i_{3}\to i_{2})$, generating only one blocking pair: $(i_{5},s_{1})$.
To conclude the proof, consider the mechanism $M$ that chooses the same matching as DA+TTC in every problem except the aforementioned one. Such matching clearly is Pareto-efficient, Pareto-dominates DA, improves more students than DA+TTC and is cardinally more stable than DA+TTC. ∎
We now discuss some implications of Theorem 3. First, our two examples reveal an interesting asymmetry: in Example 2, DA+TTC achieves the larger improvement that doubly dominates EADA, whereas in Example 3, EADA with full consent produces the improvement that double dominates DA+TTC. This observation yields the following incomparability result:
@@
EADA with full consent and DA+TTC are incomparable in terms of size of improvement and number of blocking pairs, i.e. there exist problems where DA+TTC doubly dominates EADA and vice versa.